OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points (-12,-16) and (-13,-19)

OpenStudy (nubeer):

m=(y2-y1)/(x2-x1) x1=-12,y1= -16 x2=-13,y2=-19 just plug in the values .. you will get the slope.

OpenStudy (nubeer):

@mathgirl501

OpenStudy (anonymous):

so would it be 1,13

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@nubeer

OpenStudy (anonymous):

or am i way off? I'm no good with math sorry

OpenStudy (nubeer):

you have to find the slope.. and there can't be 2 values of the slope

OpenStudy (nubeer):

well plug in the values in the formula i told above..

OpenStudy (nubeer):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

well i did that and came up with 1 and 13...is it supposed to be a fraction

OpenStudy (nubeer):

hmm well lets do again.. \[m=\frac{ -19-(-16) }{ -13-(-12) }\] tell me do u get this part? did u do it same way when you solved it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

13-12=1

OpenStudy (nubeer):

ok tell me what u got in numerator and denominator?

OpenStudy (anonymous):

thats how i got that

OpenStudy (nubeer):

infact its not 13-12 its -13-(-12) which is -13+12.. do u get it?

OpenStudy (anonymous):

oh ok I see

OpenStudy (nubeer):

ok now do it for the numerator..

OpenStudy (anonymous):

-19+16=-3

OpenStudy (nubeer):

ok good.. now tell me what fraction u have left?

OpenStudy (nubeer):

-3 in numerator and -1 in denominator..

OpenStudy (anonymous):

ok so i do subtract, i was thinking to add

OpenStudy (anonymous):

thanks

OpenStudy (nubeer):

you just have to be carful with the signs.. where to add and where to subtract so tell em what is slope?

OpenStudy (anonymous):

would it be -3

OpenStudy (nubeer):

hmm no .. look again u have -3 in numerator and -1 in denominator. so it looks like \[\frac{ -3 }{ -1 }\] minus sign will be cancelled.. so your slope will be 3..

OpenStudy (anonymous):

ok I get it now thanks for your help

OpenStudy (nubeer):

hope u understand it.. and you're welcome. ..